# Involutive Markov categories and the quantum de Finetti theorem

### 15 Dec 2023 v1

@@ 1 Introduction
A first important step is the definition of involutive Markov categories. These categories are equivalent to Parzygnat’s quantum Markov categories (Theorem 2.6), but they avoid any distinction between ‘‘even’’ and ‘‘odd’’ morphisms—which correspond to linear and antilinear maps, respectively—thus making the theory a bit cleaner. Another helpful notion introduced in this paper is that of picture, which encodes some requirements one would expect to have when dealing with quantum probability. For instance, any picture restricted to the classical objects—i.e., those encoding classical probability—becomes a Markov category. To better understand this concept, let us consider the example of C*-algebras. The associated involutive Markov category contains all unital maps, but only the completely positive ones have a well-behaved and physically relevant probabilistic interpretation. Therefore, we consider the picture given by the completely positive unital maps, and keep the involutive Markov category as the environment for diagrammatic calculus.
@@
A quantum analog of the de Finetti theorem is due to Størmer [43], and it is fundamental in quantum Bayesianism [21]. A categorical perspective on the quantum de Finetti theorem has already been considered by Staton and Summers [42], who extended Størmer’s result with parameters using categorical limits. There is, however, one major drawback: These works focus only on the minimal C*-tensor norm. The case of the maximal C*-tensor norm was considered by Hulanicki and Phelps [25], but the paper seems to be unknown. This unpopularity may be due to the fact that their proof has a minor gap; but as we show in Appendix C, their argument can easily be adapted to the case of C*-algebras with either the minimal or the maximal C*-tensor norm and proves the quantum de Finetti theorem for both. Moreover, this generalizes to the case of completely positive unital maps with an additional tensor factor (Theorem 3.68), in particular allowing for correlation and (a priori) entanglement with another system. Our result then in particular implies that there is no entanglement with the additional system, a fact whose finite-dimensional special case is well-known in quantum information theory, as it underlies a prominent separability criterion for quantum states [10, Theorem 1].
@@ Acknowledgements
We thank Tomáš Gonda, Paolo Perrone, Sam Staton and Nico Wittrock for fruitful discussions and comments. The authors acknowledge support from the Austrian Science Fund (FWF P 35992-N).
@@ 2.1 Definition and relation to quantum Markov categories
    In particular, we have $\phi^{*}=\phi\mathchoice{\,}{\,}{}{}\mathrm{id}^{*}$, and the equations are then constructed precisely such that $\overline{\phi}=\mathrm{id}^{*}\mathchoice{\,}{\,}{}{}\phi\mathchoice{\,}{\,}{}{}\mathrm{id}^{*}$. Since $\mathrm{id}^{*}\mathrm{id}^{*}=\mathrm{id}$, this also shows that an even morphism $\phi$ is self-adjoint in ${\mathsf{C}}$ in our sense if and only if $\phi\mathchoice{\,}{\,}{}{}\mathrm{id}^{*}=\mathrm{id}^{*}\mathchoice{\,}{\,}{}{}\phi$.
@@
    Theorem 2.6 tells us that quantum CD-categories and ICD-categories are different ways to talk about the same concept. This raises the question of which of these two formulations is preferable, if any. We think that ICD-categories should be preferred for defining and working with concrete examples, since only needs to consider the even morphisms. On the other hand, quantum CD-categories facilitate more general diagrammatic proofs by allowing the star *** to appear in string diagrams, as long as one is careful not to tensor an even with an odd morphism. For example, for every morphism $\phi\colon A\to B$ in an ICD-category, we can write
@@ 2.5 C*-pairs
Let $B=\mathbb{C}^{\{0,1\}}$ be the two-dimensional commutative C*-algebra whose elements are pairs of complex numbers $(b_{0},b_{1})$. Its only dense ***-subalgebra is ${}_{p}{B}=B$ itself. For any C*-pair $(A,{}_{p}{A})$, the morphisms $\phi:A\to B$ correspond to pairs of elements $(a_{0},a_{1})\in{}_{p}{A}\times{}_{p}{A}$ with $a_{0}+a_{1}=1$. Since such a pair is uniquely determined by one component, it follows that the morphisms $A\to\mathbb{C}^{\{0,1\}}$ can be identified with the elements of ${}_{p}{A}$. The self-adjoint morphisms correspond to the self-adjoint elements, while the formal opposites of positive unital maps correspond to the elements of the unit interval $[0,1]$ in ${}_{p}{A}$ [^6] As is standard for C*-algebras, the unit interval $[0,1]\subseteq{}_{p}{A}$ is the set of elements $x\in{}_{p}{A}$ such that $0\leq x\leq 1$.
While the category of ***-algebras is most naturally equipped with the algebraic tensor product as in Remarks 2.23 and 2.25, $\operatorname{C*-}{\mathsf{pr}}$ admits at least two symmetric monoidal structures, namely the minimal and the maximal tensor product of C*-algebras, which are associated to the minimal and maximal tensor norms, respectively.
@@
The categories $\operatorname{C*-}{\mathsf{pr}}_{\min}$ and $\operatorname{C*-}{\mathsf{pr}}_{\max}$ are isomorphic since we currently have no real difference between the two norms. However, this will change when we consider pictures (Definition 3.2). Indeed, the subcategories of interest will require morphisms to be bounded, and this is sensitive to the choice of the tensor C*-norm.
@@ 2.6 Classicality and compatibility
In our setting, an interesting question is which objects in an involutive Markov category can be considered ‘‘classical’’ or ‘‘commutative’’ and which ones display genuinely ‘‘quantum’’ or ‘‘noncommutative’’ behavior. With this in mind, we introduce classical objects, which, as the name suggests, are those that behave as in classical probability. Whenever an object $A$ is not classical, it is still relevant to understand how far a pair of morphisms out of $A$ is from displaying classical behavior. This gives rise to the notion of compatibility.
@@
Remark 2.39.
Instead of the classical subcategory defined above, one could alternatively consider the full subcategory on classical objects without requiring morphisms to be self-adjoint. However, the restriction to self-adjoint morphisms seems natural whenever it is possible. For example, probabilistic Gelfand duality (Proposition A.14) only considers positive unital maps, which are in particular self-adjoint. Of course, such a restriction is not possible when we have non-classical objects around, since in that case the copy morphisms are not self-adjoint and we do not want to drop such an essential piece of structure.
Notation 2.40.
@@
    This already indicates that compatibility is a notion of classicality for morphisms.
@@
In Proposition 3.12, we will turn to a slightly different version of this which actually reproduces the universal property of the maximal tensor product of C*-algebras; the present one does not, since our morphisms are not required to be bounded in any way.
@@
We first prove that the map is well-defined by showing that such $(\phi,\psi)$ is indeed deterministic. Given a pair of compatible deterministic morphisms $\phi\colon A\to B$ and $\psi\colon B\to C$, by definition of compatibility $(\phi,\psi)$ is immediately self-adjoint. The fact that it also satisfies (2.6) follows directly from (2.3).
Injectivity of the map (2.10) now immediately follows from marginalization, since $\phi$ and $\psi$ can be reconstructed from $(\phi,\psi)$ by composing with deletion on either output. For surjectivity, consider a deterministic morphism $\pi\colon A\to B\otimes C$, and take its marginals $\phi\colon A\to B$ and $\psi\colon A\to C$. As composites of deterministic morphisms, $\phi$ and $\psi$ are deterministic as well. Further we have
@@
by determinism of $\pi$ and marginalization. Finally since $\pi$ is self-adjoint, $\phi$ and $\psi$ are also compatible by Remark 2.42(iii). ∎
@@
Let $\phi,\psi$ and $\omega$ be pairwise compatible morphisms. Then $\phi$ is also compatible with $(\psi,\omega)$.
@@
The monicity assumption can now be applied to obtain the desired conclusion. ∎
@@
By autocompatibility, the right-hand side of (2.6) is invariant under swap, which means that the same is true for the left-hand side. Since $\phi$ is an epimorphism, we obtain that
$\operatorname{copy}_{B}=\operatorname{swap}\operatorname{copy}_{B},$
i.e. $B$ is classical. ∎
Combining Lemmas 2.47 and 2.49, we also obtain the following.
@@
Any morphism with classical domain is non-invasive. Whenever $\phi$ is non-invasive and self-adjoint, then also $(\phi,\mathrm{id})$ is self-adjoint. In particular, non-invasiveness of an identity morphism $\mathrm{id}_{A}$ is equivalent to the classicality of the object $A$. These facts are immediate from Remark 2.42.
@@
This notion of non-invasiveness is not to be confused with the concept of interaction-free measurement in quantum theory [46]. Such interaction-free measurements still make sense in a quantum information context, where the relevant C*-algebras are usually assumed to be matrix algebras and therefore have trivial center. In such a case, the only non-invasive morphisms are those given by discarding the original system and preparing a new state, i.e. those morphisms that factor across $I=\mathbb{C}$. On the other hand, our notion of non-invasiveness also applies to morphisms that are not measurements. [^7] Measurements are best defined as morphisms $A\to B$ for which $B$ is classical; for $B=\mathbb{C}^{n}$ in $\mathsf{CPU}_{\min/\max}$ (see Definition 3.8), this reproduces the standard notion of POVM with finitely many outcomes.
@@ 2.7 Almost sure equalities
We emphasize that $\phi\cong_{\omega}\psi$ implies $\phi\simeq_{\omega}\psi$, simply by marginalizing the left and the right output separately. The converse is not true in general, as we will see in Example 3.20 as part of a more detailed study of almost sure equalities in $\operatorname{C*-}{\mathsf{pr}}_{\min}$ and $\operatorname{C*-}{\mathsf{pr}}_{\max}$.
@@
In particular, whenever one of the two a.s. equalities holds, then $\phi\simeq_{\omega}\psi$.
Nevertheless, even in that case the fourth notion may still be different (Example 3.20).
@@ 3 Pictures, representability and the quantum de Finetti theorem
Our final topic is the quantum de Finetti theorem. This result, which we prove for both the maximal and the minimal tensor products of C*-algebras in Appendix C, is central to our theme due to its inherent connection to classical representability: we formulate it categorically as a factorization of exchangeable morphisms through the state space object (Theorem 3.63), where this formulation is rather strong in that it not only applies to exchangeable states, but to exchangeable morphisms with general input and with an extra output as additional tensor factor on the codomain.
@@ 3.1 Definition and main examples of pictures
As suggested to us by Sam Staton, a picture could alternatively be defined in terms of a strong symmetric monoidal faithful functor $U:{\mathsf{D}}\to\operatorname{SA}({\mathsf{D}})$ rather than in terms of ${\mathsf{D}}$ being a subcategory. This version may be preferred for better analogy with forgetful functors. However, we have not adopted this definition as it would clutter our notation further, while being essentially equivalent: $U$ can always be taken to be such an inclusion functor, at least modulo replacing $\operatorname{Gen}({\mathsf{D}})$ by an equivalent involutive Markov category.
@@
Another example is given by autocompatibility: if $\phi$ is an autocompatible morphism, Corollary 2.46 inductively shows that $\phi^{(n)}$, for any positive integer $n$, also belongs to the picture. In fact, this can also be reversed using Remark 2.42(iii): if $\phi^{(n)}$ is a morphism for some $n\geq 2$, then $\phi$ is autocompatible.
@@
is a Markov category. Indeed, by Property (iii) of Definition 3.2, Remark 2.42(i) implies that the copy morphisms belong to $C\ell({\mathsf{D}})$. Moreover, the same remark shows that every picture in a classical Markov category displays compatibility. This is not the case for general pictures: see the forthcoming Proposition 3.12.
@@
For the definition of this kind of maps, we refer to Definition A.8. Whenever $\mathsf{CPU}$ is considered as a subcategory of either $\operatorname{C*-}{\mathsf{pr}}_{\min}$ or $\operatorname{C*-}{\mathsf{pr}}_{\max}$, we will decorate it accordingly: $\mathsf{CPU}_{\min}$ and $\mathsf{CPU}_{\max}$, respectively. For brevity, we also use $\mathsf{CPU}_{\min/\max}$ to make statements that apply in both cases.
@@
From now on, for cleaner writing we sometimes prefer working with general elements of the containing C*-algebras rather than with elements of the dense ***-subalgebras. We will do this in cases where only positive unital maps appear, so that by Lemma A.4 all these maps extend uniquely to the containing C*-algebras.
@@
In [38], the subcategories of interest are not necessarily symmetric monoidal. In fact, the information flow axioms considered there are obtained by the Kadison–Schwarz inequality (Corollary A.11), and everything can therefore be generalized to the context of positive unital maps satisfying such an inequality. However, there does not seem to be any concrete advantage in considering such generality, since the physically realizable operations are only the completely positive unital maps, and also the mathematical literature mostly focuses on these. Furthermore, considering only symmetric monoidal subcategories provides a more understandable playground, as the reader does not have to pay attention to when tensoring is allowed. An additional motivation is given by the proof of Proposition 3.11: Property (iii) holds provided that the morphism $\psi$ is completely positive, since without this hypothesis the universal property of the maximal tensor product [5, Exercise 3.5.1] could not be applied.
@@
All morphisms are self-adjoint by Proposition A.9. Moreover, the subcategories $\mathsf{CPU}_{\min/\max}$ are symmetric monoidal since the coherence morphisms are bounded [5, Exercises 3.1.2, 3.3.1 and 3.3.2] (recall Lemma A.4), and the tensor product is respected by Proposition A.12. Moreover, the unital map $\mathbb{C}\rightsquigarrow{}_{p}{A}$ corresponding to deletion trivially belongs to our subcategories.
Concerning Property (iii), we recall Example 2.43: Given every autocompatible morphism $\phi\colon A\to B$, the completely positive map $\phi^{\mathrm{op}}$ has commutative range. In particular, $\operatorname{\operatorname{im}}(\phi^{\mathrm{op}})$ generates a commutative C*-subalgebra of $A$, and so $\phi$ admits a factorization $\phi=\phi^{\prime}\mathchoice{\,}{\,}{}{}\iota$, where $\iota^{\mathrm{op}}$ is the inclusion of a classical object. Therefore, for any morphism ${\psi}\colon A\to C$ compatible with $\phi$, the map $(\iota,\psi)^{\mathrm{op}}$ is completely positive because commutative C*-algebras are nuclear [40, Proposition 12.9], so that we can apply the universal property of the maximal tensor product [5, Exercise 3.5.1] for both monoidal structures. We conclude that $(\phi,\psi)$ is the composition of two morphisms $\phi^{\prime}\otimes\mathrm{id}$ and $(\iota,\psi)$ which are both completely positive, and so $(\phi,\psi)$ is completely positive as well. ∎
@@
One may wonder why we have not required ${\mathsf{D}}_{\mathrm{det}}={\mathsf{C}}_{\mathrm{det}}$ in the definition of picture. Indeed, between C*-algebras all ***-homomorphisms are bounded. However, this is no longer true in our setting, because we are missing the completeness of the normed spaces. A simple example is the following.
@@
It is well-known that a positive unital is always bounded (Proposition A.9). However, the same is not true for ***-homomorphisms. Equivalently, not every ***-homomorphism between pre-C*-algebras is positive.
@@
Indeed, if $\iota^{\mathrm{op}}$ was bounded for all choices of ${}_{p}{A}$ and ${}_{p}{B}$, then we would get a ***-homomorphism $\operatorname{cl}({\iota^{\mathrm{op}}})$ between C*-algebras by Lemma A.4. This $\operatorname{cl}({\iota^{\mathrm{op}}})$ would be unital and positive by Proposition A.10, so that $\mathopen{}\mathclose{{}\left\lVert\iota^{\mathrm{op}}}\right\rVert=\mathopen{}\mathclose{{}\left\lVert\operatorname{cl}({\iota^{\mathrm{op}}})}\right\rVert=1$. In particular, we would get that $\mathopen{}\mathclose{{}\left\lVert\sum_{i}a_{i}\otimes b_{i}}\right\rVert_{\max}\leq\mathopen{}\mathclose{{}\left\lVert\sum_{i}a_{i}\otimes b_{i}}\right\rVert_{\min}$ for all elements of the algebraic tensor product. Since the reverse inequality always holds, this would mean that the two norms coincide, which is generally not the case [45, Theorem 6].
A more concrete example is the following. Let us consider $A=\mathbb{C}$ and $(B,{}_{p}{B})=({C}([0,1]),\mathbb{C}[x])$. The fact that $\mathbb{C}[t]$ is a dense ***-subalgebra of $C([0,1])$ (by sending each polynomial $f$ to the continuous map $t\mapsto f(t)$) is a direct application of Stone–Weierstrass theorem. Let us now take $\phi^{\mathrm{op}}\colon\mathbb{C}[x]\rightsquigarrow\mathbb{C}$ given by evaluating polynomials at any fixed real number outside of $[0,1]$. This is a ***-homomorphism, but it cannot be bounded, because by Gelfand duality every bounded ***-homomorphism $C([0,1])\rightsquigarrow\mathbb{C}$ is given by evaluation at a point of $[0,1]$.

### 12 Jan 2024 v2

@@ 1 Introduction
A first important step is the definition of involutive Markov categories. These categories are equivalent to Parzygnat’s quantum Markov categories (Theorem 2.6), but they avoid any distinction between ‘‘even’’ and ‘‘odd’’ morphisms—which correspond to linear and antilinear maps, respectively—thus making the theory a bit cleaner. Another helpful notion introduced in this paper is that of picture, which encodes some requirements one would expect to have when dealing with quantum probability. For instance, the classical objects of any picture—i.e., those conforming to classical probability—form a Markov category. To better understand this concept, let us consider the example of C*-algebras. The associated involutive Markov category contains all unital maps, but only the completely positive ones have a well-behaved and physically relevant probabilistic interpretation. Therefore, we consider the picture given by the completely positive unital maps, and keep the involutive Markov category as the environment for diagrammatic calculus.
@@
A quantum analog of the de Finetti theorem is due to Størmer [43], and it is fundamental in quantum Bayesianism [21]. A categorical perspective on the quantum de Finetti theorem has already been considered by Staton and Summers [42], who extended Størmer’s result with parameters using categorical limits. There is, however, one major drawback: These works focus only on the minimal C*-tensor norm. The case of the maximal C*-tensor norm was considered by Hulanicki and Phelps [25], but the paper seems to be unknown. Their argument, which holds for both the minimal and the maximal C*-tensor norm, is presented in Appendix C. Moreover, this generalizes to the case of completely positive unital maps with an additional tensor factor (Theorem 3.69), in particular allowing for correlation and (a priori) entanglement with another system. Our result then in particular implies that there is no entanglement with the additional system, a fact whose finite-dimensional special case is well-known in quantum information theory, as it underlies a prominent separability criterion for quantum states [10, Theorem 1].
@@ Acknowledgements
We thank Tomáš Gonda, Paolo Perrone, Sam Staton, Ned Summers and Nico Wittrock for fruitful discussions and comments. The authors acknowledge support from the Austrian Science Fund (FWF P 35992-N).
@@ 2.1 Definition and relation to quantum Markov categories
    In particular, we have $\phi^{*}=\phi\mathchoice{\,}{\,}{}{}\mathrm{id}^{*}$, and the equations are then constructed precisely such that $\overline{\phi}=\mathrm{id}^{*}\mathchoice{\,}{\,}{}{}\phi\mathchoice{\,}{\,}{}{}\mathrm{id}^{*}$. Since $\mathrm{id}^{*}\mathrm{id}^{*}=\mathrm{id}$, this also shows that an even morphism $\phi$ is self-adjoint in ${\mathsf{C}}$ in our sense if and only if $\phi\mathchoice{\,}{\,}{}{}\mathrm{id}^{*}=\mathrm{id}^{*}\mathchoice{\,}{\,}{}{}\phi$, i.e. $\phi$ is ***-preserving.
@@
    Theorem 2.6 tells us that quantum CD-categories and ICD-categories are different ways to talk about the same concept. This raises the question of which of these two formulations is preferable, if any. We think that ICD-categories should be preferred for defining and working with concrete examples, since one only needs to consider even morphisms. On the other hand, quantum CD-categories facilitate more general diagrammatic proofs by allowing the star *** to appear in string diagrams, as long as one is careful not to tensor an even with an odd morphism. For example, for every morphism $\phi\colon A\to B$ in an ICD-category, we can write
@@ 2.5 C*-pairs
Let $B=\mathbb{C}^{\{0,1\}}$ be the two-dimensional commutative C*-algebra whose elements are pairs of complex numbers $(b_{0},b_{1})$, multiplication is defined componentwise, and the unit is $(1,1)$. Its only dense ***-subalgebra is ${}_{p}{B}=B$ itself. For any C*-pair $(A,{}_{p}{A})$, the morphisms $\phi:A\to B$ correspond to pairs of elements $(a_{0},a_{1})\in{}_{p}{A}\times{}_{p}{A}$ with $a_{0}+a_{1}=1$—such elements are the images of $(1,0)$ and $(0,1)$ respectively. Since such a pair is uniquely determined by one component, it follows that the morphisms $A\to\mathbb{C}^{\{0,1\}}$ can be identified with the elements of ${}_{p}{A}$. The self-adjoint morphisms correspond to the self-adjoint elements, while the formal opposites of positive unital maps correspond to the elements of the unit interval $[0,1]$ in ${}_{p}{A}$ [^6] As is standard for C*-algebras, the unit interval $[0,1]\subseteq{}_{p}{A}$ is the set of elements $x\in{}_{p}{A}$ such that $0\leq x\leq 1$.
While the category of ***-algebras is most naturally equipped with the algebraic tensor product as in Remark 2.23, $\operatorname{C*-}{\mathsf{pr}}$ admits at least two symmetric monoidal structures, namely the minimal and the maximal tensor product of C*-algebras, which are associated to the minimal and maximal tensor norms, respectively.
@@
The categories $\operatorname{C*-}{\mathsf{pr}}_{\min}$ and $\operatorname{C*-}{\mathsf{pr}}_{\max}$ are isomorphic, since the norms do not currently play any role. However, this will change when we consider pictures (Definition 3.2). Indeed, the subcategories of interest will require morphisms to be bounded, and this is sensitive to the choice of the tensor C*-norm.
@@ 2.6 Classicality and compatibility
In our setting, an interesting question is which objects in an involutive Markov category can be considered ‘‘classical’’ or ‘‘commutative’’ and which ones display genuinely ‘‘quantum’’ or ‘‘noncommutative’’ behavior. With this in mind, we introduce classical objects, which, as the name suggests, are those that behave as in classical probability. Whenever an object $A$ is not classical, it is still relevant to understand when a pair of morphisms out of $A$ displays ‘‘classical’’ behavior. This gives rise to the notion of compatibility.
@@
Notation 2.39.
@@
In Proposition 3.12, we will turn to a slightly different version of this result, which actually reproduces the universal property of the maximal tensor product of C*-algebras. The present one does not, since here the morphisms are not required to be bounded in any way.
@@
We first prove that the map is well-defined by showing that such $(\phi,\psi)$ is indeed deterministic. Given a pair of compatible deterministic morphisms $\phi\colon A\to B$ and $\psi\colon B\to C$, by definition of compatibility $(\phi,\psi)$ is immediately self-adjoint. It is also deterministic, since the copy morphisms respect the symmetric monoidal structure (2.3).
Injectivity of the map (2.10) now immediately follows from marginalization, since $\phi$ and $\psi$ can be reconstructed from $(\phi,\psi)$ by composing with deletion on either output. For surjectivity, consider a deterministic morphism $\pi\colon A\to B\otimes C$, and take its marginals
@@
As composites of deterministic morphisms, $\phi$ and $\psi$ are deterministic as well. Further we have
[…]
by determinism of $\pi$. Finally since $\pi$ is self-adjoint, $\phi$ and $\psi$ are also compatible by Remark 2.41(iii). ∎
@@
Let $\phi,\psi$, and $\omega$ be pairwise compatible morphisms. Then $\phi$ is also compatible with $(\psi,\omega)$.
@@
The monicness assumption can now be applied to obtain the desired conclusion. ∎
@@
By autocompatibility, $(\phi,\phi)$ is invariant under swap. Since $\phi$ is deterministic, we infer that
[…]
The epicness of $\phi$ implies that the same holds when $\phi$ is dropped; hence $B$ is classical. ∎
Combining Lemmas 2.46 and 2.48, we also obtain the following.
@@
From Remark 2.41 we deduce the following facts:
  1. (i)
    Any morphism with classical domain is non-invasive.
  2. (ii)
    Whenever $\phi$ is non-invasive and self-adjoint, then also $(\phi,\mathrm{id})$ is self-adjoint. In particular, the non-invasiveness of the identity morphism $\mathrm{id}_{A}$ of $A$ is equivalent to the classicality of the object $A$.
@@
This notion of non-invasiveness is not to be confused with the concept of interaction-free measurement in quantum theory [46]. While a general definition of what counts as ‘‘interaction-free’’ does not seem to exist, we can see as follows that our notion of non-invasiveness is quite different. If $A$ is a matrix algebra (of matrix size $>1$), then its centre is trivial, and therefore the only non-invasive morphisms $A\to B$ are those that factor across $I=\mathbb{C}$. This is a physically uninteresting operation, as it amounts to discarding the original system and preparing a new state. But the case of matrix algebras is the one that most quantum information literature focuses on, including the literature on interaction-free measurements, which can still be interesting and nontrivial despite trivial center.
On the other hand, our notion of non-invasiveness also applies to morphisms that are not measurements. [^7] Measurements can be defined as morphisms $A\to B$ for which $B$ is classical; for $B=\mathbb{C}^{n}$ in $\mathsf{CPU}_{\min/\max}$ (see Definition 3.8), this reproduces the standard notion of POVM with finitely many outcomes.
@@ 2.7 Almost sure equalities
We emphasize that $\phi\cong_{\omega}\psi$ implies $\phi\simeq_{\omega}\psi$, simply by marginalizing the left and the right output of (2.12) separately. The converse is not true in general, as we will see in Example 3.20 as part of a more detailed study of almost sure equalities in $\operatorname{C*-}{\mathsf{pr}}_{\min}$ and $\operatorname{C*-}{\mathsf{pr}}_{\max}$.
@@
In particular, whenever one of the two a.s. equalities holds, then $\phi\simeq_{\omega}\psi$ also holds.
Nevertheless, even in that case the symmetric a.s. equality may still be different (Example 3.20).
@@ 3 Pictures, representability and the quantum de Finetti theorem
Our final topic is the quantum de Finetti theorem. This result, which we prove for both the maximal and the minimal tensor products of C*-algebras in Appendix C, is central to our topic because of its inherent connection to classical representability, as it is categorically formulated as a factorization of exchangeable morphisms through the state space object (Theorem 3.64). This formulation is rather strong in that it not only applies to exchangeable states, but to exchangeable morphisms with general input and with an extra output as an additional tensor factor on the codomain.
@@ 3.1 Definition and main examples of pictures
As suggested to us by Sam Staton, a picture could alternatively be defined in terms of a strong symmetric monoidal faithful functor $U\colon{\mathsf{D}}\to\operatorname{SA}({\mathsf{C}})$ rather than in terms of ${\mathsf{D}}$ being a subcategory. This version may be preferred for better analogy with forgetful functors. However, we have not adopted this definition as it would clutter our notation further, while being essentially equivalent: $U$ can always be taken to be such an inclusion functor, at least modulo replacing $\operatorname{Gen}({\mathsf{D}})$ by an equivalent involutive Markov category.
@@
Another example is given by autocompatibility: if $\phi$ is an autocompatible morphism, an inductive argument combining Corollary 2.45 and Property (iii) shows that $\phi^{(n)}$, for any positive integer $n$, also belongs to the picture. In fact, this can also be reversed using Remark 2.41(iii): if $\phi^{(n)}$ is a morphism for some $n\geq 2$, then $\phi$ is autocompatible.
@@
is a Markov category. Indeed, by Property (iii) of Definition 3.2, Remark 2.41(i) implies that the copy morphisms belong to $C\ell({\mathsf{D}})$.
@@
For the definition of such maps, we refer to Definition A.8. Whenever $\mathsf{CPU}$ is considered as a subcategory of either $\operatorname{C*-}{\mathsf{pr}}_{\min}$ or $\operatorname{C*-}{\mathsf{pr}}_{\max}$, we will decorate it accordingly: $\mathsf{CPU}_{\min}$ and $\mathsf{CPU}_{\max}$, respectively. For brevity, we also use $\mathsf{CPU}_{\min/\max}$ to make statements that apply to both cases.
@@
From now on, for cleaner writing we sometimes prefer working with general elements of the containing C*-algebras $A$ rather than with elements of the dense ***-subalgebras ${}_{p}{A}$. We employ this only in contexts where all maps involved are positive unital, since such maps have a unique extension to the containing C*-algebras (see Lemma A.4).
@@
In [38], the focus is on subcategories that are not necessarily symmetric monoidal. In fact, the information flow axioms considered there are obtained by the Kadison–Schwarz inequality (Corollary A.11), and all results can therefore be stated in the context of positive unital maps satisfying such an inequality. However, there does not seem to be any concrete advantage in considering such generality, since the physically realizable operations are only the completely positive unital maps, and also the mathematical literature mostly focuses on these. Furthermore, considering only symmetric monoidal subcategories provides a more understandable playground, as the reader does not have to pay attention to when tensoring is allowed. An additional motivation is given by the proof of Proposition 3.11: Property (iii) holds provided that the morphisms considered are completely positive, since without this hypothesis the universal property of the maximal tensor product [5, Exercise 3.5.1] could not be applied.
@@
All morphisms are self-adjoint by Proposition A.9. Moreover, the subcategories $\mathsf{CPU}_{\min/\max}$ are symmetric monoidal since the coherence morphisms are bounded [5, Exercises 3.1.2, 3.3.1 and 3.3.2] (recall Lemma A.4), and the tensor product is respected by Proposition A.12. Moreover, the unique unital map $\mathbb{C}\rightsquigarrow{}_{p}{A}$ corresponding to deletion is completely positive.
Concerning Property (iii), we recall Example 2.42: For any autocompatible morphism $\phi\colon A\to B$, the completely positive map $\phi^{\mathrm{op}}$ has commutative range. In particular, the range $\operatorname{\operatorname{im}}(\phi^{\mathrm{op}})$ generates a commutative C*-subalgebra $R\subseteq A$, and so $\phi$ admits a factorization $\phi=\phi^{\prime}\mathchoice{\,}{\,}{}{}\iota$, where $\iota^{\mathrm{op}}\colon R\rightsquigarrow A$ is the inclusion. Since any morphism ${\psi}\colon A\to C$ compatible with $\phi$ is also compatible with $\iota$, the map $(\iota,\psi)^{\mathrm{op}}$ is therefore completely positive because commutative C*-algebras are nuclear [40, Proposition 12.9], so that we can apply the universal property of the maximal tensor product [5, Exercise 3.5.1] for both monoidal structures. We conclude that $(\phi,\psi)$ is the composition of two morphisms $\phi^{\prime}\otimes\mathrm{id}$ and $(\iota,\psi)$ which are both completely positive, and so $(\phi,\psi)$ is completely positive as well. ∎
@@
One may wonder why we have not required ${\mathsf{D}}_{\mathrm{det}}=\operatorname{Gen}({\mathsf{D}})_{\mathrm{det}}$ in the definition of picture. Indeed, between C*-algebras all ***-homomorphisms are bounded. However, this is no longer true for C*-pairs, because we are missing the completeness of the normed spaces. A simple example is the following.
@@
It is well-known that a positive unital map is always bounded (Proposition A.9). However, the same is not true for ***-homomorphisms. Equivalently, not every ***-homomorphism between pre-C*-algebras is positive.
@@
Indeed, if $\iota^{\mathrm{op}}$ was bounded for all choices of ${}_{p}{A}$ and ${}_{p}{B}$, then we would get a ***-homomorphism $\operatorname{cl}({\iota^{\mathrm{op}}})\colon A\otimes_{\min}B\to A\otimes_{\max}B$ by Lemma A.4. This $\operatorname{cl}({\iota^{\mathrm{op}}})$ would be unital and positive by Proposition A.10, so that $\mathopen{}\mathclose{{}\left\lVert\iota^{\mathrm{op}}}\right\rVert=\mathopen{}\mathclose{{}\left\lVert\operatorname{cl}({\iota^{\mathrm{op}}})}\right\rVert=1$. In particular, we would get that $\mathopen{}\mathclose{{}\left\lVert\sum_{i}a_{i}\otimes b_{i}}\right\rVert_{\max}\leq\mathopen{}\mathclose{{}\left\lVert\sum_{i}a_{i}\otimes b_{i}}\right\rVert_{\min}$ for all elements of the algebraic tensor product. Since the reverse inequality always holds, this would mean that the two norms coincide, which is generally not the case [45, Theorem 6].
A more concrete example is the following. Let us consider $A=\mathbb{C}$ and $(B,{}_{p}{B})=({C}([0,1]),\mathbb{C}[x])$, where $C([0,1])$ is the C*-algebra of complex-valued continuous maps on the unit interval $[0,1]$, and $\mathbb{C}[x]$ is included in $C([0,1])$ by identifying each polynomial $f\in\mathbb{C}[x]$ with the associated function $x\mapsto f(x)$. By the Stone–Weierstrass theorem, such an identification makes $\mathbb{C}[x]$ a dense ***-subalgebra of $C([0,1])$. Let us now take $\phi^{\mathrm{op}}\colon\mathbb{C}[x]\rightsquigarrow\mathbb{C}$ given by evaluating polynomials at any fixed real number outside of $[0,1]$. This is a ***-homomorphism, but it cannot be bounded, because by Gelfand duality every bounded ***-homomorphism $C([0,1])\rightsquigarrow\mathbb{C}$ is given by evaluation at a point of $[0,1]$.